Calculate the pH of a buffer solution

formed by dissolving 0.0350 mole

HCH3COO and 0.550 mole of NaCH3COO

in 0.950 L solution. Ka of HCH3COO = 1.8 x

-5 10 .


1
Expert's answer
2021-02-04T07:03:22-0500

Molar concentration of CH3COOH = (0.035/0 950)=0.036mol/L

Molar concentration of CH3COONa = (0.55/0.950)=0 578mol/L

Pka= -log(Ka)= 4.74

This is an acedic buffer solution

pH= pKa+ log(CH3COONa/CH3COOH)

=5.94